# Fusion of Movement and Naïve Predictions for Point Forecasting in Univariate Random Walks

### v2

@@ Abstract
Traditional methods for point forecasting in univariate random walks often fail to surpass naïve benchmarks due to data unpredictability. This study introduces a novel forecasting method that fuses movement prediction (binary classification) with naïve forecasts for accurate one-step-ahead point forecasting. The method’s efficacy is demonstrated through theoretical analysis, simulations, and real-world data experiments. It reliably exceeds naïve forecasts with movement prediction accuracies as low as 0.55, outperforming baseline models like ARIMA, linear regression, MLP, and LSTM networks in forecasting the S&P 500 index and Bitcoin prices. This method is particularly advantageous when accurate point predictions are challenging but accurate movement predictions are attainable, translating movement predictions into point forecasts in random walk contexts.
@@ 1 Introduction
Another key characteristic of a random walk is that future values are considered unpredictable based on past values alone, as predicting a random walk is akin to predicting a series of random events Fama (1995); Zhang (1999). Despite significant efforts, including complex modeling using both machine learning and deep learning algorithms, along with various levels of data fusion for data augmentation Thakkar and Chaudhari (2021), forecasting models often struggle to effectively recognize meaningful patterns in random walk series and typically underperform naïve forecasting, which assumes that the future value of the series will be the same as the most recent value Ellwanger and Snudden (2023); Hewamalage et al. (2023). Recent studies have proposed various time-frequency decomposition methods that have been shown to outperform naïve forecasts in random walk forecasting Hewamalage et al. (2023); Zhang et al. (2023). However, these methods often include future information in the training data, leading to data leakage and a higher forecasting accuracy Hewamalage et al. (2023); Wu et al. (2022); Zhang et al. (2015). When data leakage is explicitly avoided during data preprocessing, the prediction accuracy often falls below that of the naïve forecast Hewamalage et al. (2023). This superiority of naïve forecasts underscores the significant challenge of random walk point forecasting Moosa and Burns (2014).
@@
The theory of prediction involves constructing mathematical models that attempt to describe real phenomena and using these models to forecast future outcomes Frigg and Hartmann (2018). However, most regression methods for time series forecasting presuppose a mapping from historical values to future values, which conflicts with the fundamental concept of random walk, where the future value depends solely on the current value plus a random increment and higher-order autocorrelations rapidly diminish toward zero. This misalignment in the underlying prediction mechanisms can significantly contribute to the suboptimal performance of these methods, particularly when the assumed patterns and dependencies do not exist in the data with the true random walk nature.
@@
This study addresses the challenge of random walk point forecasting from a novel perspective, focusing on translating high-accuracy movement predictions into point forecasts that can surpass naïve forecasts, the potential of which remains underexplored and underappreciated. By introducing a sign representing the movement direction of the future value, the random error term in the traditional definition of random walk, which is independent and identically distributed (i.i.d.), can be expressed as a positive i.i.d. random error multiplied by this sign. With regard to prediction, the sign can be estimated via movement prediction, and the estimation of the positive error for the future point can be obtained empirically based on the estimation of movement prediction accuracy and the residual of the naïve forecast on the historical data. Therefore, from this variant definition of random walk, the one-step-ahead point forecasting of a univariate random walk can theoretically be derived from the fusion of movement and naïve predictions. The movement prediction integrated in this fusion method can be achieved by using various machine learning and deep learning algorithms, which are advancing rapidly.
@@
  • By leveraging binary classification, which is typically considered unrelated to regression tasks, point forecasting of random walks that surpass naïve forecasts becomes possible. In areas such as financial time series forecasting, predicting the direction of movement (a binary classification task) can be easier than forecasting the exact value of a target variable Taylor (2008). Although complex feature engineering using external information can enhance both movement and point predictions, the latter generally remains less accurate than naïve forecasts. The proposed method shows that random walk point forecasting can be accurate to a certain degree when the movement prediction accuracy reaches a specific level.
  • This study highlights directions for future improvements in random walk forecasting. Rather than developing end-to-end regression models, we should instead focus on enhancing movement prediction accuracy. Accurate binary classification is equivalent to accurate regression in univariate random walk forecasting.
@@ 2.1 Point forecasting of random walks
Accurate point forecasting for random walks is challenging. In financial markets, the behavior of financial instruments often resembles a random walk, presenting few, if any, discernible patterns for analysts to exploit Fama (1995). Nonetheless, several forecasting methods have been developed to address this challenge.
@@
Furthermore, convolutional recurrent neural networks (CRNNs), which integrate the spatial pattern recognition capabilities of convolutional neural networks (CNNs) with the temporal analysis strengths of RNNs, can enhance the landscape of time series point forecasting Shi et al. (2017). This hybrid approach has been effectively applied to price forecasting, illustrating its potential for discerning spatiotemporal patterns in complex time series data Tsantekidis et al. (2020). Moreover, transformer-based models, such as Vision Transformer (ViT) and Autoformer, have been adapted to process financial time series due to their ability to capture intricate patterns and dependencies over extended periods Malibari et al. (2021); Wu et al. (2021). ViT, originally designed for image classification, utilizes self-attention mechanisms to manage global interactions within sequential data, making it adaptable for financial time series analysis. Autoformer, designed specifically for time series forecasting, incorporates decomposition blocks and autocorrelation mechanisms to enhance its effectiveness in modeling long-term temporal dependencies. These properties make transformer-based approaches increasingly useful for understanding and predicting the dynamics of time series.
@@ 2.2 Movement prediction of random walks
Various machine learning and deep learning models that utilize a range of advanced techniques have been proposed for movement prediction. Classical and machine learning models employ technical indicators, price data, and fundamental data for movement prediction Huynh et al. (2017); Raza (2017); Sezer et al. (2017), whereas neural networks integrated with selective trading strategies and ensemble networks have been shown to be effective in movement prediction Das et al. (2018); Yang et al. (2017). Additionally, LSTM and gated recurrent unit (GRU) networks, incorporating technical indicators such as moving average convergence divergence (MACD), demonstrate their utility in handling movement estimation for stock markets Nelson et al. (2017); Troiano et al. (2018).
@@
Notably, point forecasting and movement prediction were explored independently. The potential of translating accurate binary classification to superior point forecasting is largely underappreciated.
@@ 3 Method
This section discusses the details of the proposed fusion method, by which movement predictions and naïve forecasts are integrated for point forecasting. Derived from a variant definition of random walk, the fusion method was introduced, followed by a theoretical analysis of the fusion results, which emphasizes how improvements in movement prediction accuracy can influence fusion outcomes. Aligned with the aim of this study, which is to demonstrate the effectiveness of the proposed fusion method for random walk point forecasting, this section omits details on movement prediction models and only discusses the fusion process at the decision level. Any model capable of producing meaningful movement predictions can be integrated to improve point forecasts through the proposed fusion method.
@@ 3.1 Fusion of movement and naïve predictions
where ${{M}_{t}}\in\{-1,1\}$ is a sign factor that indicates the changing direction of the data point value at time step $t$ compared with the data point value at time step $t-1$ and $\epsilon_{t}^{+}$ is a positive i.i.d. random error term that represents the magnitude of the increment between these two adjacent points. As these positive increments are uncorrelated, they are theoretically unpredictable. Nonetheless, $\epsilon_{t}^{+}$ can further be expressed as a multiplication of two positive values:
@@
where $\bar{\epsilon}$ is the mean of the absolute residuals of the naïve forecast calculated from the historical data, which is a constant positive value. The term ${{c}_{t}}$, which is positive and i.i.d., is introduced as a weight that controls the magnitude of the increment at time step $t$. Suppose a simplified scenario in which the increment between two adjacent points is constant for each future time step; then, ${{c}_{t}}$ should be set to a fixed value in advance. Consequently, the forecast of future value, ${{\hat{y}}_{t}}$, can be expressed as follows:
@@
where ${{\hat{M}}_{t}}$ is the movement prediction, $c$ is a fixed scalar, and $c\in R$. Given that $\bar{\epsilon}$ is derived from the naïve forecast and that the optimal value of $c$ depends only on the movement prediction accuracy (proved in the following theoretical analysis), the point forecast, based on (4), can be regarded as a fusion of movement and naïve predictions. Suppose movement prediction ${{\hat{M}}_{t}}$ is provided by a binary classification model, and the value of $c$ is given; then, the point forecast for the whole test set can be calculated using Algorithm 1.
@@
$P$: Array of forecasted values for the test set
@@ 2. Generate naïve forecasts for the $Y_{\text{train}}$:
$Y_{\text{naive}}\leftarrow Y_{\text{train}}[1:\text{end}-1]$
@@ 3. Determine the residuals from naïve forecasts:
$R\leftarrow Y_{\text{train}}[2:\text{end}]-Y_{\text{naive}}$
@@ 5. Initialize the prediction array and set the initial forecast:
6. Forecast each point in $Y_{\text{test}}$ by integrating the movement predictions:
@@
     $\Delta\leftarrow c\cdot\bar{\epsilon}\times M[i]$ // Increment calculation using mean absolute residual and direction
@@ 3.2 Theoretical analysis of the fusion results
This subsection provides a theoretical analysis of the performance of the fusion method by comparing the MSE, the fundamental metric for quantifying the accuracy of point forecasting, between the fusion and naïve predictions on the test set. Assuming that both the training and test sets are sufficiently large, with respective counts of $M$ and $N$ data points, and that the movement prediction results in ${{n}_{correct}}$ correct directions and ${{n}_{incorrect}}$ incorrect directions for the test set, the MSE of the fusion results on the test set can be expressed as follows:
@@
$MSE^{\text{naive}}$ $=\frac{1}{N}\left(\sum_{t\in\text{testset}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (6)
@@
$MSE^{\text{naive}}-MSE^{\text{fusion}}$ $=\frac{1}{N}\left(\sum_{t\in\text{correct}}(\epsilon_{t}^{+})^{2}+\sum_{t\in\text{incorrect}}(\epsilon_{t}^{+})^{2}\right)$ (10)
@@
Based on the above analysis, interestingly, the optimal value of $c$ depends on $ACC^{\text{movement}}$. Given that when $AC{{C}^{movement}}\in(0.5,1)$, the movement prediction is meaningful, we can obtain the corresponding optimal value of $c$, which falls within the range between 0 and 1. If we set a large value for $c$, the threshold of $ACC^{\text{movement}}$ for improvement from the naïve forecast must also be high. Based on (11), when the $ACC^{\text{movement}}$ is only 0.51, setting a small value for $c$, such as 0.01, can theoretically ensure that the fusion forecast outperforms the naïve forecast in terms of the MSE. The value assigned to the weight $c$ for the fusion prediction represents a trade-off between the likelihood of outperforming the naïve forecast and the magnitude of the improvement. When $c$ is fixed, an increase in $ACC^{\text{movement}}$ directly increases the discrepancy in the MSE between the naïve and fusion forecasts. In a real forecasting scenario where $ACC^{\text{movement}}$ is unknown, we can use the movement prediction accuracy on the training set or validation set as an estimation of $ACC^{\text{movement}}$ and set $c$ based on (11). Because the RMSE is the root of the MSE, obtaining the requirement for $RMS{{E}^{fusion}}<RMS{{E}^{naive}}$ is equivalent to obtaining that for $MS{{E}^{fusion}}<MS{{E}^{naive}}$.
According to (4), (10), and (11), point forecasting based on the fusion method will be affected only by $ACC^{\text{movement}}$, as $c$ is set based on an estimation of $ACC^{\text{movement}}$, and there is no uncertainty for either the naïve forecast or $\bar{\epsilon}$ once the historical data are determined. This finding implies that accurate movement prediction is equivalent to accurate point forecasting in univariate random walks.
@@ 4 Experimental results
To evaluate the effectiveness of the proposed method for random walk forecasting, this study employed a two-pronged experimental approach, providing a holistic view of its performance across different conditions. Initially, the method was tested using simulated data to understand its performance under theoretical conditions. Subsequently, the method was applied to real-world data to assess its practical utility and effectiveness in actual forecasting scenarios.
The experiments were conducted on a cloud-based Jupyter notebook service, which provided adaptable virtual machines with CPUs for simulated experiments and L4 GPUs for training baseline models such as MLP and LSTM. The software environment included Python 3.8 with core libraries such as TensorFlow 2.3 and NumPy 1.19. All the methods in the experiments were evaluated using the MAE and MSE. The formulas for these metrics are as follows:

### v3

@@ Abstract
Traditional methods for point forecasting in univariate random walks often fail to surpass naïve benchmarks due to data unpredictability. This study introduces a novel forecasting method that fuses movement prediction (binary classification) with naïve forecasts for accurate one-step-ahead point forecasting in univariate random walks. The method’s efficacy is demonstrated through theoretical analysis, simulations, and real-world data experiments. It reliably outperforms naïve forecasts with moderate movement prediction accuracies, such as 0.55, and is superior to baseline models such as the ARIMA, linear regression, MLP, and LSTM networks in forecasting the S&P 500 index and Bitcoin prices. This method is particularly advantageous when accurate point predictions are challenging but accurate movement predictions are attainable, translating movement predictions into point forecasts in random walk contexts.
@@ 1 Introduction
Another key characteristic of a random walk is that future values are considered unpredictable based on past values alone, as predicting a random walk is akin to predicting a series of random events Fama (1995); Zhang (1999). Despite significant efforts, including complex modeling using both machine learning and deep learning algorithms, along with various levels of data fusion for data augmentation Thakkar and Chaudhari (2021), forecasting models often struggle to effectively recognize meaningful patterns in random walk series and typically underperform naïve forecasting, which assumes that the future value of the series will be the same as the most recent value Ellwanger and Snudden (2023); Hewamalage et al. (2023). Recent studies have proposed various time-frequency decomposition methods that have been shown to outperform naïve forecasts in random walk forecasting Hewamalage et al. (2023); Zhang et al. (2023). However, these methods often include future information in the training data, leading to data leakage and a higher forecasting accuracy Hewamalage et al. (2023); Wu et al. (2022); Zhang et al. (2015). When data leakage is explicitly avoided during data preprocessing, the prediction accuracy often falls below that of the naïve forecast Hewamalage et al. (2023). This superiority of naïve forecasts underscores the significant challenge of point forecasting in random walks Moosa and Burns (2014).
@@
Prediction tasks in different areas often involve constructing mathematical models that attempt to describe real phenomena and using these models to forecast future outcomes Frigg and Hartmann (2018). However, most regression methods implemented for random walk forecasting presuppose a mapping from historical values to future values, which does not align with the concept of random walks, where the future value depends solely on the current value with a random increment and higher-order autocorrelations rapidly diminish toward zero. This misalignment in the underlying prediction mechanisms can significantly contribute to the suboptimal performance of these methods, particularly when the assumed patterns and dependencies do not exist in the data with the true random walk nature.
@@
This study addresses the challenge of random walk point forecasting from a novel perspective, focusing on translating movement predictions into point forecasts, the potential of which remains less investigated. By introducing a sign representing the movement direction of the future value, the random error term in the definition of random walks, which is independent and identically distributed (i.i.d.), can be expressed as a positive i.i.d. random error multiplied by this sign. With regard to prediction, the sign can be estimated via movement prediction, and the estimation of the positive error for the future point can be obtained empirically based on both the estimation of movement prediction accuracy and the residual of the naïve forecast on the historical data. Therefore, from this variant definition of random walks, the one-step-ahead point forecasting of a univariate random walk can theoretically be derived from the fusion of movement and naïve predictions. More importantly, this fusion method does not require a loss function for optimization. Instead, it can be optimized by estimating movement prediction accuracy on the validation set, avoiding rolling back to a traditional regression method that often faces bottlenecks in random walk forecasting. Furthermore, movement prediction can be achieved using various machine learning and deep learning algorithms, which are advancing rapidly.
@@
  • By leveraging binary classification, which is typically considered unrelated to regression tasks, point forecasting of random walks that surpass naïve forecasts becomes possible. In financial time series forecasting, complex feature engineering using external information can enhance movement and point predictions. However, point predictions generally remain less accurate than naïve forecasts, and predicting the direction of movement (a binary classification task) is often easier than forecasting the exact value of a target variable Taylor (2008). The proposed method shows that random walk point forecasting can be predicted to a certain degree by incorporating movement prediction with a certain accuracy.
  • This study highlights directions for future improvements in random walk forecasting. Rather than developing end-to-end regression models, we could instead focus on enhancing movement prediction accuracy. The findings of this study show that accurate binary classification is equivalent to accurate regression in univariate random walk forecasting.
@@ 2.1 Point forecasting of random walks
Accurate point forecasting for random walks has always been challenging. In financial markets, the behavior of financial instruments often resembles a random walk, presenting few, if any, discernible patterns for analysts to exploit Fama (1995). Nonetheless, several forecasting methods have been developed to address this challenge.
@@
Furthermore, convolutional recurrent neural networks (CRNNs), which integrate the spatial pattern recognition capabilities of convolutional neural networks (CNNs) with the temporal analysis strengths of RNNs, have been effectively applied to price forecasting, illustrating the potential for discerning spatiotemporal patterns in complex time series data Shi et al. (2017); Tsantekidis et al. (2020). Moreover, transformer-based models, such as Vision Transformer (ViT) and Autoformer, have been adapted to process financial time series due to their ability to capture intricate patterns and dependencies over extended periods Malibari et al. (2021); Wu et al. (2021). ViT, originally designed for image classification, utilizes self-attention mechanisms to manage global interactions within sequential data, making it adaptable for financial time series analysis. Autoformer, designed specifically for time series forecasting, incorporates decomposition blocks and autocorrelation mechanisms to enhance its effectiveness in modeling long-term temporal dependencies. These properties make transformer-based approaches increasingly useful for understanding and predicting the dynamics of time series.
@@ 2.2 Movement prediction of random walks
Various machine learning and deep learning models that utilize a range of advanced techniques have been proposed for movement prediction. Classical and machine learning models employ technical indicators, price data, and fundamental data for movement prediction Huynh et al. (2017); Raza (2017); Sezer et al. (2017), whereas neural networks integrated with selective trading strategies have been shown to be effective in movement prediction Das et al. (2018); Yang et al. (2017). Additionally, LSTM and gated recurrent unit (GRU) networks, incorporating technical indicators such as moving average convergence divergence (MACD), demonstrate their utility in handling movement estimation for stock markets Nelson et al. (2017); Troiano et al. (2018).
@@
Notably, point forecasting and movement prediction were explored independently. The potential of translating accurate binary classification to superior point forecasting has been insufficiently studied.
@@ 3 Method
This section discusses the details of the proposed fusion method, by which movement predictions and naïve forecasts are integrated for point forecasting. Derived from a variant definition of random walks, the fusion method was introduced, followed by a theoretical analysis of the fusion results, which emphasizes how improvements in movement prediction accuracy can influence fusion outcomes. Aligned with the aim of this study, which is to demonstrate the effectiveness of the proposed fusion method for random walk point forecasting, this section omits details on movement prediction models and only discusses the fusion process at the decision level. Any model capable of producing meaningful movement predictions can be integrated to improve point forecasts through the proposed fusion method.
@@ 3.1 Fusion of movement and naïve predictions
where ${{M}_{t}}\in\{-1,1\}$ is a sign factor that indicates the changing direction of the point value at time step $t$ compared with the point value at time step $t-1$ and $\epsilon_{t}^{+}$ is a positive i.i.d. random error term that represents the magnitude of the increment between these two adjacent points. As these positive increments are uncorrelated, they are theoretically unpredictable. Nonetheless, $\epsilon_{t}^{+}$ can further be expressed as a multiplication of two positive values:
@@
where $\bar{\epsilon}$ is the mean of the absolute residuals of the naïve forecast calculated from the historical data, which is a constant positive value. The term ${{c}_{t}}$, which is positive and i.i.d., is introduced as a weight that controls the magnitude of the increment at time step $t$. Suppose a simplified scenario in which the increment between two adjacent points is constant for future time steps; then, ${{c}_{t}}$ can be set to a fixed value in advance. Consequently, the forecast of future value, ${{\hat{y}}_{t}}$, can be expressed as follows:
@@
where ${{\hat{M}}_{t}}$ is the movement prediction, $c$ is a fixed scalar, and $c\in R$. The point forecast, based on (4), can be regarded as a fusion of movement and naïve predictions. Given that ${y}_{t-1}$ and $\bar{\epsilon}$ are determined from the historical data and ${{\hat{M}}_{t}}$ is provided by movement prediction, optimizing the fusion result is equivalent to finding the optimal value of $c$. Suppose movement prediction ${{\hat{M}}_{t}}$ is provided by a binary classification model, and the value of $c$ is given; then, the fusion result for the whole test set can be calculated using Algorithm 1.
@@
$P$: Array of fusion results for the test set
@@ 2. Generate naïve forecasts for the $Y_{\text{train}}$:
$Y_{\text{na\"{\i}ve}}\leftarrow Y_{\text{train}}[1:\text{end}-1]$
@@ 3. Determine the residuals from naïve forecasts:
$R\leftarrow Y_{\text{train}}[2:\text{end}]-Y_{\text{na\"{\i}ve}}$
@@ 5. Initialize the prediction array and set the initial forecast:
6. Each point in $Y_{\text{test}}$ is forecasted by integrating the movement predictions:
@@
     $\Delta\leftarrow c\cdot\bar{\epsilon}\times M[i]$ // Increment calculation using the mean absolute residual and direction
@@ 3.2 Theoretical analysis of the fusion results
This subsection presents a theoretical analysis of the performance of the fusion method by comparing the MSE, the fundamental metric for quantifying the accuracy of point forecasting, between the fusion and naïve predictions on the test set. This theoretical analysis also provides a method to determine the optimal value of c to achieve the best fusion result. Assuming that both the training and test sets are sufficiently large, with respective counts of $M$ and $N$ data points, and that the movement prediction results in ${{n}_{correct}}$ correct directions and ${{n}_{incorrect}}$ incorrect directions for the test set, the MSE of the fusion results on the test set can be expressed as follows:
@@
$MSE^{\text{na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{testset}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (6)
@@
$MSE^{\text{na\"{\i}ve}}-MSE^{\text{fusion}}$ $=\frac{1}{N}\left(\sum_{t\in\text{correct}}(\epsilon_{t}^{+})^{2}+\sum_{t\in\text{incorrect}}(\epsilon_{t}^{+})^{2}\right)$ (10)
@@
Based on the above analysis, interestingly, the optimal value of $c$ depends only on $ACC^{\text{movement}}$. Given that when $AC{{C}^{movement}}\in(0.5,1)$, the movement prediction is meaningful, we can obtain the corresponding optimal value of $c$, which falls within the range between 0 and 1. If we set a large value for $c$, the threshold of $ACC^{\text{movement}}$ for improvement from the naïve forecast must also be high. Based on (11), when the $ACC^{\text{movement}}$ is only 0.51, setting a small value for $c$, such as 0.01, can theoretically ensure that the fusion forecast outperforms the naïve forecast in terms of the MSE. The value assigned to the weight $c$ for the fusion prediction represents a trade-off between the likelihood of outperforming the naïve forecast and the magnitude of the improvement. When $c$ is fixed, an increase in $ACC^{\text{movement}}$ directly increases the discrepancy in the MSE between the naïve and fusion forecasts. In a real forecasting scenario where $ACC^{\text{movement}}$ is unknown, we can use the movement prediction accuracy on the training set or validation set as an estimation of $ACC^{\text{movement}}$ and set $c$ based on (11). Because the RMSE is the root of the MSE, obtaining the requirement for $RMS{{E}^{fusion}}<RMS{{E}^{naive}}$ is equivalent to obtaining that for $MS{{E}^{fusion}}<MS{{E}^{naive}}$.
According to (4), (10), and (11), point forecasting based on the fusion method will be affected only by $ACC^{\text{movement}}$, and the fusion result is optimized by setting the value of $c$ based on an estimation of $ACC^{\text{movement}}$. This finding implies that accurate movement prediction can be translated to accurate point forecasting in univariate random walks.
@@ 4 Experimental results
To evaluate the effectiveness of the proposed method for random walk forecasting, this study employed a two-pronged experimental approach, providing a holistic view of its performance across different conditions. Initially, the method was tested using simulated data to understand its performance under different levels of $ACC^{\text{movement}}$. Subsequently, the method was applied to real-world data to assess its practical utility and effectiveness in actual forecasting scenarios.
The experiments were conducted on a cloud-based Jupyter notebook service, which provided adaptable virtual machines with CPUs for simulated experiments and L4 GPUs for training baseline models such as MLP and LSTM. The software environment included Python 3.8 with core libraries such as TensorFlow 2.3 and NumPy 1.19. All the methods in the experiments were evaluated using the MSE and MAE. The formulas for these metrics are as follows:
$MSE=\frac{1}{N}\sum\limits_{i=1}^{N}{{{({{y}_{i}}-{{{\hat{y}}}_{i}})}^{2}}},$ (12)